Solve the system by graphing both equations on the same axes. Give the ordered pair for the point of intersection.
x-2y=6
x+2y=2

Answer by waynest(281)

x -2y = 6 subtract these two equations
-(x +2y = 2)
-4y = 4
y = -1
x -2(-1) = 6
x + 2 = 6
x = 4
(4, -1)